OpenStudy (anonymous):

Simplify. 70 + (–62) A. 132 B. 8 C. –8 D. –132

OpenStudy (anonymous):

@ali2x2 @dreamstoogood

OpenStudy (anonymous):

@ccieux

OpenStudy (lynfran):

whats 70-62=?

OpenStudy (zzr0ck3r):

\(a+(-b) = a-b\)

OpenStudy (anonymous):

I got B. 8 What did you get?